40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat May 21 06:19:46 2011, kiwamu, Update, LSC, DRMI trial : no luck 
    Reply  Mon May 23 00:36:51 2011, kiwamu, Update, LSC, DRMI trial : sucess 
       Reply  Mon May 23 12:27:26 2011, kiwamu, Update, LSC, DRMI trial : details 
Message ID: 4760     Entry time: Mon May 23 12:27:26 2011     In reply to: 4759
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: DRMI trial : details 

(PRMI locking with slightly misaligned SRM)

 First I tried locking PRC and MICH with a little bit misaligned SRM. This condition allowed me to search for a good signal port for SRC.

In this locking, REFL11_I was used to lock PRC and AS55_Q was used for MICH. This is the same scheme as the current PRMI locking.

Since the alignment of SRM was close to the good alignment, I expected to see fringes from SRC in some signal ports (i.e. REFL55, POY55 and so on).

Sometimes a fringe of SRC disturbed AS55_Q and broke the MICH locking, so I had to carefully misalign SRM so that the SRC fringes are small enough to maintain the lock of MICH.

 

(Looking for a good signal port for SRC)

 After I locked the PRMI with slightly misaligned SRM, I started looking for a good signal port for SRC.

At the beginning I tried finding a good SRC port by shaking SRM at 100 Hz and looked at the power spectra of all the available LSC signals.

I was expecting to see a 100 Hz peak in the spectra, but this technique didn't work well because SRC wasn't within the linear range and hence didn't produce linear signals.

So I didn't see any strong signals at 100 Hz and finally gave up this technique.

Then I started looking for a PDH-like signal in time series and immediately found AS55_I showed large PDH-like signals.

So I started using the AS55_I for the SRC locking and eventually succeeded.
 

 

(Two tips for the DRMI locking)

During the locking of DRMI, I found two tips that made the locking quite smooth.

 - Triggered locking

   Since every LSC signal ports showed large signals from PRC somehow, feeding back the signals made the suspensions crazy.

   So I used triggered locking for the PRC and MICH locking to avoid unwanted kicks on BS and PRM.

   If  the DC of REFL goes above a certain level, the control of  PRC starts. Also if the DC of AS goes below a certain level the control of MICH starts.

  These triggers make the lock smoother.

 - Do not use resonant gain filters

  This is really a stupid tip. When I was trying to lock MICH, the lock became quite difficult for some reasons.

  It looked there was an oscillation at 3 Hz every time the MICH control started. It turned out that a 3 Hz resonant gain filter had been making it difficult.

  All the resonant gain filters should be off when a lock acquisition is taken place.

Quote from #4759

Eventually the DRMI was locked.

More details will be reported in the morning.

 

ELOG V3.1.3-